OpenStudy (anonymous):

If (82)p = 84, what is the value of p?

OpenStudy (freckles):

you can find p by dividing both sides by what is next to p

OpenStudy (anonymous):

so 82 divived by 84?

OpenStudy (freckles):

the other way around

OpenStudy (freckles):

remember you divided both sides by 82

OpenStudy (anonymous):

oh ok thanks

OpenStudy (freckles):

\[\frac{\cancel{82} p}{\cancel{82}}=\frac{84}{82} \\ p=\frac{84}{82}\] but this can be reduced

OpenStudy (anonymous):

k that should help

OpenStudy (freckles):

k cool stuff